OpenStudy (mathmath333):

How many natural numbers not more than 4300 can be formed with the digits 0,1,2,3,4 (if repetitions are allowed)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ How many natural numbers not more than 4300}\hspace{.33em}\\~\\ & \normalsize \text{ can be formed with the digits 0,1,2,3,4}\hspace{.33em}\\~\\ & \normalsize \text{ (if repetitions are allowed)}\hspace{.33em}\\~\\ & a.)\ 574 \hspace{.33em}\\~\\ & b.)\ 570 \hspace{.33em}\\~\\ & c.)\ 575 \hspace{.33em}\\~\\ & d.)\ 569 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmate):

Using regex notation and abbreviating [01234] as [] Note: [01234] means choose one digit out of {0,1,2,3,4} [0123][][][] = 4*5^3 -[0][0][0][0] = -1 [4][012][[][] = 1*3*5^2 [4][0][0][0] = 1 total = 500-1+75+1=575 (if natural number defined as {1,2,3,4.....}]

OpenStudy (mathmate):

*[4][3][0][0] = 1

OpenStudy (mathmath333):

i m not getting this notation

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & 5^{4}-1-(2\times 5^{2} -1)=575 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmate):

Yes, you can do that too!